# Hadron-quark hybrid model, modular transformation and Roberge-Weiss transition

### v1

@@ Abstract
In the framework of modular transformations, we reformulate the recently proposed hadron-quark hybrid model when the imaginary baryonic chemical potential is introduced. As a result, we can consider the torus, which is characterized by the complex number densities of baryons (antibaryons) and quarks (antiquarks). We apply this model to analyze the Roberge-Weiss transition. It is shown that the torus vanishes at the baryonic chemical potential where the Roberge-Weiss transition appears because the number density of baryons (antibaryons) is not linearly independent of the number density of quarks (antiquarks). When the temperature $T$ is lower than the Roberge-Weiss transition temperature $T_{\rm RW}$, the torus shrinks smoothly to the one-dimensional object at the Roberge-Weiss transition point, but the discontinuity does not appear. On the other hand, the discontinuity of the geometrical object appears when $T>T_{\rm RW}$. We also calculate the modulus of the torus and transform it into the fundamental region. The transformed moduli are symmetric below $T_{\rm RW}$, but the symmetry is broken above $T_{\rm RW}$.
@@ I Introduction
Determination of the phase diagram of quantum chromodynamics (QCD) is an important subject not only in nuclear and particle physics but also in cosmology and astrophysics; see, e.g., Ref. Fukushima:2010bq and references therein. However, when the baryon number chemical potential $\mu$ is finite and real, the fist principle calculation, that is, the lattice QCD (LQCD) simulation, is not feasible due to the infamous sign problem; see Refs. deForcrand:2010ys; Nagata:2021bru; *Nagata:2021ugx as an example. To circumvent the sign problem, several methods are proposed and investigated, although, at present, these methods are not complete and we do not have adequate information on the equation of state (EoS) at finite real baryon density.
One possible way to avoid the sign problem is to use the LQCD results with the imaginary baryon number chemical potential; see Refs. deForcrand:2002ci; D'Elia:2002gd; D'Elia:2004at; Chen:2004tb; D'Elia:2009qz as an example. When the baryon number chemical potential $\mu$ is pure imaginary, there is no sign problem. One can perform LQCD simulations at finite pure imaginary $\mu$, and then make an analytic continuation from the quantities at imaginary $\mu$ to those at real $\mu$. Alternatively, one may determine the unknown parameters of an effective model of QCD using the LQCD results at imaginary $\mu$. After determining the parameters, the model calculations can be performed at real $\mu$ Sakai:2009dv. One can also construct the canonical partition function with fixed baryon number from the grand canonical partition function with the pure imaginary chemical potential  Roberge:1986mm. In principle, the grand canonical partition function with the real chemical potential can be constructed from the canonical ones. However, at present, these approaches are far from perfection.
QCD at imaginary $\mu$ itself has very interesting properties and may provide us with an important insight into the physical QCD at real $\mu$. The grand canonical partition function $Z(\theta_{\mathrm{Q}})$ with pure imaginary quark chemical potential ($\mu_{\mathrm{Q}}={\mu/3}=i\theta_{\mathrm{Q}}T$) has the Roberge-Weiss (RW) periodicity Roberge:1986mm as
@@
where $T$ is the temperature and $\theta_{\mathrm{Q}}\in\mathbb{R}$. This periodicity is the remnant of the $\mathbb{Z}_{3}$-symmetry of pure gluon theory. At low temperature, $Z(\theta_{\mathrm{Q}})$ is a smooth function of $\theta_{\mathrm{Q}}$. However, at high temperature above the RW temperature $T_{\rm RW}$, it has a singularity at $\theta_{\mathrm{Q}}=(2k+1)\pi/3$ where $k\in\mathbb{Z}$. This singularity is called the RW transition Roberge:1986mm. Figure 1 shows a schematic phase diagram of the RW transitions. The order of the endpoint of the RW transition line, which is called the RW endpoint, has not yet been determined definitively. $T_{\rm RW}$ for 2+1 flavor QCD is estimated as about $200$ MeV by LQCD simulations Bonati:2016pwz; Cuteri:2022vwk; Bonati:2018fvg. The RW transition is well described by quark models such as the Polyakov-loop extended Nambu–Jona-Lasinio (PNJL) model Sakai:2008py; Kouno:2009bm. However, such a quark model fails to reproduce the EoS of QCD matter at low temperature.
@@
The RW transition is expected to have a strong relation to the deconfinement transition. A construction of the hybrid model which can reproduce the low- and high-temperature properties of LQCD results as well as the RW transition is desirable. In this paper, we reformulate the recently proposed hadron-quark hybrid model  Kouno:2023ygw in the framework of the modular transformation when the pure imaginary baryonic chemical potential is introduced. Using the model, we qualitatively analyze the RW transition. It is shown that we can consider the torus which is characterized by the quark and baryon densities of the hybrid model and is invariant under the modular transformation, but the torus vanishes at $\theta=(2k+1)\pi$ ($\theta_{\mathrm{Q}}=(2k+1)\pi/3$). When $T<T_{\rm RW}$, the torus shrinks smoothly to the one-dimensional object at $\theta=(2k+1)\pi$, but the discontinuity does not appear. On the other hand, the discontinuity of the geometrical object appears when $T>T_{\rm RW}$.
@@ II Roberge-Weiss periodicity and transition
Hence, we obtain the RW periodicity (1). Dynamical quarks break the $\mathbb{Z}_{3}$-symmetry but the RW periodicity appears as a remnant of the $\mathbb{Z}_{3}$-symmetry Roberge:1986mm. In the following, we concentrate our discussions on the region $0\leq\theta_{\rm Q}\leq{2\pi\over{3}}~{}(0\leq\theta\leq 2\pi)$, since the properties of the RW transition in the other regions are similar to those of this region.
@@
See Fig. 2 and, also see, e.g., Fig. 7 in Ref. Sakai:2009dv and the references therein.
@@
This relation is also well explained by the effective model of QCD, such as the Polyakov-loop extended Nambu–Jona-Lasinio (PNJL) model  Fukushima:2003fw; Ratti:2005jh; Ghosh:2006qh; Megias:2004hj; Roessner:2006xn; Kashiwa:2007hw; Sakai:2010rp; Sakai:2009dv; Sakai:2008py; Kouno:2009bm. In the PNJL model, the $\theta_{\rm Q}$-dependence appears only in the form of $\Phi\,e^{i\theta_{\rm Q}}=|\Phi|\,e^{i(\phi+\theta_{\rm Q})}$. Hence, at low temperature, $\phi$ tends to cancel the effects of $\theta$ and the approximate relation $\phi=-\theta$ holds. At high temperature the $\theta_{\rm Q}$-independent Polyakov-loop potential that originates in the pure gluon sector of QCD has $\mathbb{Z}_{3}$-symmetric three minima with ϕ=italic-$\phi=$0, $-{2\pi\over{3}}$ and $-{4\pi\over{3}}$, respectively. Due to the effects of the quark part that breaks the $\mathbb{Z}_{3}$ symmetry explicitly, one of these minima becomes the vacuum of the system. When $0\leq\theta<{\pi\over{3}}$, $\phi=0$, but $\phi$ jumps from 0 to $-{2\pi\over{3}}$ at $\theta_{\rm Q}={\pi\over{3}}~{}(\theta=\pi)$, that is, at the RW transition point.
It is well known that the PNJL model can reproduce the other several important features of QCD; for example, see Ref. Kashiwa:2019ihm as a review. However, at low temperature, the PNJL model fails to reproduce the EoS of the LQCD simulation. Hence, a hybrid model that simultaneously includes the quark and hadron degrees of freedom is needed.
@@ III Modular transformation
in the complex plane. That is, $z_{1}$ and $z_{2}$ are the bases of this lattice. We can consider the torus defined by the quotient space $\mathbb{C}/\Lambda$. $\tau$ is called a ”modulus” of the torus. The modular transformation (20) is a base conversion, but this conversion does not change the lattice and, hence, the structures of the torus.
The modulus of the torus has ambiguity, since the modular transformation does not change the structure of the torus. However, it is known that $\tau$ can be transformed into the following fundamental region $D_{\rm up}$ by the series of S and T transformations (hence, by the modular transformation), when ${\rm Im}(\tau)>0$;
@@ IV Hadron-quark hybrid model
First, we briefly review the hybrid model recently proposed in Ref Kouno:2023ygw. Suppose $n_{\rm b}$ and $n_{\rm a}$ are the number densities of point-like baryons and antibaryons, respectively. We can effectively include the repulsive interaction among baryons (anitbaryons) in the model by considering the excluded volume effects (EVE). For the EVE, see, e.g., Ref. Kouno:2023ygw and the references therein. The modified number density $\tilde{n}_{\rm b}$ ($\tilde{n}_{\rm a}$) of baryons (antibaryons) is given by
@@
where $v$ is the volume of a baryon (an antibaryon). The net baryon number density is given by $n_{\rm b}-n_{\rm a}$. Other thermodynamic quantities such as the pressure and the energy density can be obtained by using thermodynamic relations. However, when $n_{\rm b}~{}(n_{\rm a})\to\infty$, $\tilde{n}_{\rm b}$ ($\tilde{n}_{\rm a}$) approaches the constant value $1/v$, the EoS of baryon matter becomes very hard and the speed of sound exceeds 1  Kouno:2023ygw. The causality is easily violated. Hence, the $T$ and/or $\mu$ dependence of $v$ is very important.
@@
It should be remarked that the relation $v=1/n_{\rm q}$ is not valid in the limit $T,\mu\to 0$ where $n_{\rm q}$ also approaches 00. It is known that nucleon has a finite volume $v_{0}={4\pi\over{3}}r_{0}^{3}$ with $r_{0}\sim 0.8$ fm in vacuum. In Ref. Kouno:2023ygw $v$ which interpolates $1/n_{\rm q}$ and $v_{0}$ is used. However, in this paper, we restrict our discussion to the region where $T$ is above or just below $T_{\rm RW}$. Hence, we use the simple relation $v=1/n_{\rm q}$ in this paper.
@@
(2) Equation (47) does not change the lattice and torus. Hence, only the information of the untransformed quantities $n_{\rm q}$ and $n_{\rm b}$ is needed to analyze the properties of the lattice and torus.
@@ V.1 Vanishing of the torus
In particular, it seems that Eq. (55) holds exactly when $\theta=0,\pi$ in LQCD simulation. See, e.g., Fig. 7 in Ref Sakai:2009dv. and references therein. The phase of the modulus $\tau$ is given by
@@
At $\theta=\pi$, $n_{\rm b}$ is antiparallel to $n_{\rm q}$, and, hence, it is not linearly independent with $n_{\rm q}$. There, the torus shrinks smoothly to a one-dimensional object. Although there is no discontinuity of $\Delta\phi_{\mathrm{b}-q}$ below $T_{\rm RW}$, in this sense, the point $\theta=\pi$ is special. Of course, $n_{\rm b}$ is parallel to $n_{\rm q}$ at $\theta=0$, but this fact is trivial since $\theta$ itself vanishes.
@@ V.2 Calculation of moduli in the fundamental region
Figure 4 shows the moduli $\tau$ and $\tau_{\rm f}$ when $T<T_{\rm RW}$ and $|\tau|=0.5$. $\tau$ and $\tau_{\rm f}$ are symmetric with respect to the line ${\rm Re}(\tau)=0$. There is a left-right symmetry. Although the torus vanishes at $\theta=\pi$, $\tau$ is continuous there. On the other hand, $\tau_{\rm f}$ shows divergent behavior there. These divergent behaviors of $\tau_{\rm f}$ may be related to the disappearance of the torus.
Fig. 4: $\tau$ and $\tau_{\rm f}$ when $T<T_{\rm RW}$ and $|\tau|=0.5$. The dot-dashed line and the dots are $\tau$ and $\tau_{\rm f}$, respectively. $\tau(\theta\to 0)=(0.5,0)$, $\tau(\theta\to\pi\pm 0)=(-0.5,0)$ and $\tau(\theta\to 2\pi)=(0.5,0)$, while $\tau_{\rm f}(\theta\to 0)$, $\tau_{\rm f}(\theta\to\pi\pm 0)$ and $\tau_{\rm f}(\theta\to 2\pi)$ show divergent behaviors.
Fig. 5: $\tau$ and $\tau_{\rm f}$ when $T<T_{\rm RW}$ and $|\tau|=0.95$. The dot-dashed line and the dots are $\tau$ and $\tau_{\rm f}$, respectively. $\tau(\theta\to 0)=(0.95,0)$, $\tau(\theta\to\pi\pm 0)=(-0.95,0)$ and $\tau(\theta\to 2\pi)=(0.95,0)$ while $\tau_{\rm f}(\theta\to 0)$, $\tau_{\rm f}(\theta\to\pi\pm 0)$ and $\tau_{\rm f}(\theta\to 2\pi)$ show divergent behaviors.
@@
Fig. 8: $\tau$ and $\tau_{\rm f}$ when $T>T_{\rm RW}$ and $|\tau|=1.05$. The dot-dashed line and the dots are $\tau$ and $\tau_{\rm f}$, respectively. $\tau(\theta\to 0)=(1.05,0)$, $\tau(\theta\to\pi-0)=(-1.05/2,1.05\sqrt{3}/2)$, $\tau(\theta\to\pi+0)=(-1.05/2,-1.05\sqrt{3}/2)$, $\tau(\theta\to 2\pi)=(1.05,0)$, $\tau_{\rm f}(\theta\to\pi-0)=(0.475,0.909)$ and $\tau_{\rm f}(\theta\to\pi+0)=(0.475,-0.909)$ while $\tau_{\rm f}(\theta\to 0)$ and $\tau_{\rm f}(\theta\to 2\pi)$ show divergent behavior.
Fig. 9: $\tau$ and $\tau_{\rm f}$ when $T>T_{\rm RW}$ and $|\tau|=2$. The dot-dashed line and the dots are $\tau$ and $\tau_{\rm f}$, respectively. $\tau(\theta\to 0)=(2.0,0)$, $\tau(\theta\to\pi-0)=(-1.0,\sqrt{3})$, $\tau(\theta\to\pi+0)=(-1.0,-\sqrt{3})$, $\tau_{\rm f}(\theta\to\pi-0)=(0,\sqrt{3})$ and $\tau_{\rm f}(\theta\to\pi+0)=(0,-\sqrt{3})$ while $\tau_{\rm f}(\theta\to 0)$ and $\tau_{\rm f}(\theta\to 2\pi)$ show divergent behaviors.
@@
Figure 7 shows the moduli $\tau$ and $\tau_{\rm f}$ when $T\to T_{\rm RW}+0$ and $|\tau|=1$. $\tau$ is not symmetric with respect to the line ${\rm Re}(\tau)=0$, but $\tau_{\rm f}$ is, since the left boundary of the fundamental region can be transformed into the right one by the T transformation and vice versa. Also note that $\tau_{\rm f}$ in Fig. 7 resembles that in Fig. 6, but the value of $\tau_{\rm f}(\theta\to\pi)$ is different. In this sense, the symmetry is broken even for $\tau_{\rm f}$ in Fig. 7. Also note that $\tau=(-0.5,\pm\sqrt{3}/2)$ is a fixed point of the modular transformation ${\cal{ST}}$. This transformation generates a discrete subgroup of the modular group, namely, ${\cal{Z}}_{3}$ group, since $({\cal{ST}})^{3}$ is an identity transformation. It is known that this property is related to the residual ${\cal{Z}}_{3}$-symmetry in modular flavor symmetry Kikuchi:2022geu.
@@
In Fig. 10, we summarize the image of the RW transition based on the torus deformations. When $T<T_{\rm RW}$, the torus shrinks to the one-dimensional object at $\theta=\pi$ smoothly, and then the object transforms into another torus $\mathrm{Im}(\tau)$ has an opposite sign, but discontinuity does not appear. Since the modular transformation cannot change the sign of $\mathrm{Im}(\tau)$, the torus with a positive $\mathrm{Im}(\tau)$ cannot be smoothly transformed into the one with a negative $\mathrm{Im}(\tau)$ without shrinking of the torus. When, $T>T_{\rm RW}$, the torus changes discontinuously at $\theta=\pi$. The structure of the torus with modulus $\tau$ at high $T$ is similar to that with modulus $-1/\tau$ at low $T$ but the left-right symmetry is broken at high $T$.
@@
Since $s^{2}-1/4$ is real positive constant which does not depend on $\theta$, $\tau_{\mathrm{h}}$ forms a circle. The circle vanishes in the limit $|\tau|\to 0$ or $|\tau|\to\infty$. The center of the circle is given by

### v4

@@ Abstract
In the framework of modular transformations, we reformulate the recently proposed hadron-quark hybrid model when the imaginary baryonic chemical potential is introduced. In this case, the number density of the hybrid model is obtained by the modular transformation of the complex number densities of the baryons (antibaryons) and the quarks (antiquarks). We can regard these number densities as the basis in the complex plane. As a result, we can consider the torus which is characterized by the basis. Since the complex structure of the torus is invariant under the modular transformation, we can extract the topological property of the hadron-quark system using the untransformed baryon (antibaryon) and the quark (antiquark) number densities. We apply this model to analyze the Roberge-Weiss transition. It is shown that the torus vanishes at the baryonic chemical potential where the Roberge-Weiss transition appears because the number density of baryons (antibaryons) is not linearly independent of the number density of quarks (antiquarks). When the temperature $T$ is lower than the Roberge-Weiss transition temperature $T_{\rm RW}$, the torus shrinks smoothly to the one-dimensional object at the Roberge-Weiss transition point, but the discontinuity does not appear. On the other hand, the discontinuity of the geometrical object appears when $T>T_{\rm RW}$. We also calculate the modulus of the torus and transform it into the fundamental region. The transformed moduli are symmetric below $T_{\rm RW}$, but the symmetry is broken above $T_{\rm RW}$.
@@ I Introduction
Determination of the phase diagram of quantum chromodynamics (QCD) is an important subject not only in nuclear and particle physics but also in cosmology and astrophysics; see, e.g., Ref. Fukushima and Hatsuda (2011) and references therein. However, when the baryon number chemical potential $\mu$ is finite and real, the fist principle calculation, that is, the lattice QCD (LQCD) simulation, is not feasible due to the infamous sign problem; see Refs. de Forcrand (2009); Nagata (2020); *Nagata:2021ugx as an example. To circumvent the sign problem, several methods are proposed and investigated, although, at present, these methods are not complete and we do not have adequate information on the equation of state (EoS) at finite real baryon density.
One possible way to avoid the sign problem is to use the LQCD results with the imaginary baryon number chemical potential; see Refs. de Forcrand and Philipsen (2002); D’Elia and Lombardo (2003, 2004); Chen and Luo (2005); D’Elia and Sanfilippo (2009) as an example. When the baryon number chemical potential $\mu$ is pure imaginary, there is no sign problem. One can perform LQCD simulations at finite pure imaginary $\mu$, and then make an analytic continuation from the quantities at imaginary $\mu$ to those at real $\mu$. Alternatively, one may determine the unknown parameters of an effective model of QCD using the LQCD results at imaginary $\mu$. After determining the parameters, the model calculations can be performed at real $\mu$ Sakai et al. (2009). One can also construct the canonical partition function with fixed baryon number from the grand canonical partition function with the pure imaginary chemical potential  Roberge and Weiss (1986). In principle, the grand canonical partition function with the real chemical potential can be constructed from the canonical ones. However, at present, these approaches are far from perfection.
QCD at imaginary $\mu$ itself has very interesting properties and may provide us with an important insight into the physical QCD at real $\mu$. The grand canonical partition function $Z(\theta_{\mathrm{Q}})$ with pure imaginary quark chemical potential ($\mu_{\mathrm{Q}}={\mu/3}=i\theta_{\mathrm{Q}}T$) has the Roberge-Weiss (RW) periodicity Roberge and Weiss (1986) as
@@
where $T$ is the temperature and $\theta_{\mathrm{Q}}\in\mathbb{R}$. This periodicity is the remnant of the $\mathbb{Z}_{3}$-symmetry of pure gluon theory. At low temperature, $Z(\theta_{\mathrm{Q}})$ is a smooth function of $\theta_{\mathrm{Q}}$. However, at high temperature above the RW temperature $T_{\rm RW}$, it has a singularity at $\theta_{\mathrm{Q}}=(2k+1)\pi/3$ where $k\in\mathbb{Z}$. This singularity is called the RW transition Roberge and Weiss (1986). Figure 1 shows a schematic phase diagram of the RW transitions. The order of the endpoint of the RW transition line, which is called the RW endpoint, has not yet been determined definitively. $T_{\rm RW}$ for 2+1 flavor QCD is estimated as about $200$ MeV by LQCD simulations Bonati et al. (2016); Cuteri et al. (2022); Bonati et al. (2019). The RW transition is well described by quark models such as the Polyakov-loop extended Nambu–Jona-Lasinio (PNJL) model Sakai et al. (2008); Kouno et al. (2009). However, such a quark model fails to reproduce the EoS of QCD matter at low temperature.
@@
The RW transition is expected to have a strong relation to the deconfinement transition. A construction of the hybrid model which can reproduce the low- and high-temperature properties of LQCD results as well as the RW transition is desirable. In this paper, we reformulate the recently proposed hadron-quark hybrid model  Kouno and Kashiwa (2024) in the framework of the modular transformation when the pure imaginary baryonic chemical potential is introduced. The baryonic number denisty $\tilde{n}_{\mathrm{b}}$ of the hybrid model is given by
$\tilde{n}_{\rm b}={n_{\rm b}\over{1+(n_{\rm b}/n_{\rm q})}}.$ (3)
where $n_{\mathrm{b}}$ and $n_{\mathrm{q}}$ are the number density of baryons and a one-third of the number density of quarks, respectively. Similarly, the antibaryonic number density of the hybrid model is given using the number densities of antibaryons and antiquarks. When $n_{\mathrm{q}}$ is nonzero, Eq. (3) can be rewritten as
$\tau^{\prime}\equiv{\tilde{n}_{\rm b}\over{n_{\mathrm{q}}}}={\tau\over{\tau+1}},$ (4)
where $\tau\equiv n_{\mathrm{b}}/n_{\mathrm{q}}$. At the imaginary baryonic chemical potential, $\tau$ is a complex number in general. Hence, Eq. (4) is a kind of modular transformations. For detail description of the modular transformation, see Sec. III.
We regard $n_{\mathrm{b}}$ and $n_{\mathrm{q}}$ as the basis in the complex plane. As a result, we can consider the lattice and the torus, which are characterized by the basis, or equivalently, their ratio $\tau$ which is called modulus of the torus. For a brief review of the lattice and the torus, see appendix A. Since the complex structure of the torus is invariant under the modular transformation (4), we can extract the topological property of the hadron-quark system using the untransformed baryon (antibaryon) and the quark (antiquark) number densities. Using the torus, we qualitatively analyze the RW transition. The torus vanish if $n_{\mathrm{b}}$ and $n_{\mathrm{q}}$ is not linearly independent. In fact, when $T<T_{\rm RW}$, the torus shrinks smoothly to the one-dimensional object at $\theta=(2k+1)\pi~{}(\theta_{\mathrm{Q}}=(2k+1)\pi/3)$, but the discontinuity does not appear. On the other hand, the discontinuity of the geometrical object appears there when $T>T_{\rm RW}$.
@@ II Roberge-Weiss periodicity and transition
Hence, we obtain the RW periodicity (1). Dynamical quarks break the $\mathbb{Z}_{3}$-symmetry but the RW periodicity appears as a remnant of the $\mathbb{Z}_{3}$-symmetry Roberge and Weiss (1986). In the following, we concentrate our discussions on the region $0\leq\theta_{\rm Q}\leq{2\pi\over{3}}~{}(0\leq\theta\leq 2\pi)$, since the properties of the RW transition in the other regions are similar to those of this region.
@@
See Fig. 2 and, also see, e.g., Fig. 7 in Ref. Sakai et al. (2009) and the references therein.
@@
This relation is also well explained by the effective model of QCD, such as the Polyakov-loop extended Nambu–Jona-Lasinio (PNJL) model  Dumitru and Pisarski (2002); Fukushima (2004); Ratti et al. (2006); Ghosh et al. (2006); Megias et al. (2006); Rößner et al. (2007); Kashiwa et al. (2008); Sakai et al. (2010, 2009, 2008); Kouno et al. (2009). In the PNJL model, the $\theta_{\rm Q}$-dependence appears only in the form of $\Phi\,e^{i\theta_{\rm Q}}=|\Phi|\,e^{i(\phi+\theta_{\rm Q})}$. Hence, at low temperature, $\phi$ tends to cancel the effects of $\theta$ and the approximate relation $\phi=-\theta$ holds. At high temperature the $\theta_{\rm Q}$-independent Polyakov-loop potential that originates in the pure gluon sector of QCD has $\mathbb{Z}_{3}$-symmetric three minima with ϕ=italic-$\phi=$0, $-{2\pi\over{3}}$ and $-{4\pi\over{3}}$, respectively. Due to the effects of the quark part that breaks the $\mathbb{Z}_{3}$ symmetry explicitly, one of these minima becomes the vacuum of the system. When $0\leq\theta<{\pi\over{3}}$, $\phi=0$, but $\phi$ jumps from 0 to $-{2\pi\over{3}}$ at $\theta_{\rm Q}={\pi\over{3}}~{}(\theta=\pi)$, that is, at the RW transition point.
It is well known that the PNJL model can reproduce the other several important features of QCD; for example, see Ref. Kashiwa (2019) as a review. However, at low temperature, the PNJL model fails to reproduce the EoS of the LQCD simulation. Hence, a hybrid model that simultaneously includes the quark and hadron degrees of freedom is needed.
@@ III Modular transformation
in the complex plane. That is, $z_{1}$ and $z_{2}$ are the basis of this lattice. We can consider the torus defined by the quotient space $\mathbb{C}/\Lambda$. $\tau$ is called a ”modulus” of the torus. See Appendix A. The modular transformation (22) is a base conversion, but this conversion does not change the similarity of lattice and, hence, the complex structures of the torus.
The modulus of the torus has ambiguity, since the modular transformation does not change the complex structure of the torus. However, it is known that $\tau$ can be transformed into the following fundamental region $D_{\rm up}$ by the series of S and T transformations (hence, by the modular transformation), when ${\rm Im}(\tau)>0$;
@@ IV Hadron-quark hybrid model
First, we briefly review the hybrid model recently proposed in Ref Kouno and Kashiwa (2024). Suppose $n_{\rm b}$ and $n_{\rm a}$ are the number densities of point-like baryons and antibaryons, respectively. We can effectively include the repulsive interaction among baryons (antibaryons) in the model by considering the excluded volume effects (EVE). For the EVE, see, e.g., Ref. Kouno and Kashiwa (2024) and the references therein. The modified number density $\tilde{n}_{\rm b}$ ($\tilde{n}_{\rm a}$) of baryons (antibaryons) is given by
@@
where $v$ is the volume of a baryon (an antibaryon). The net baryon number density is given by $n_{\rm b}-n_{\rm a}$. Other thermodynamic quantities such as the pressure and the energy density can be obtained by using thermodynamic relations. However, when $n_{\rm b}~{}(n_{\rm a})\to\infty$, $\tilde{n}_{\rm b}$ ($\tilde{n}_{\rm a}$) approaches the constant value $1/v$, the EoS of baryon matter becomes very hard and the speed of sound exceeds 1  Kouno and Kashiwa (2024). The causality is easily violated. Hence, the $T$ and/or $\mu$ dependence of $v$ is very important.
@@
It should be remarked that the relation $v=1/n_{\rm q}$ is not valid in the limit $T,\mu\to 0$ where $n_{\rm q}$ also approaches 00. It is known that nucleon has a finite volume $v_{0}={4\pi\over{3}}r_{0}^{3}$ with $r_{0}\sim 0.8$ fm in vacuum. In Ref. Kouno and Kashiwa (2024) $v$ which interpolates $1/n_{\rm q}$ and $v_{0}$ is used. However, in this paper, we restrict our discussion to the region where $T$ is above or just below $T_{\rm RW}$. Hence, we use the simple relation $v=1/n_{\rm q}$ in this paper.
@@
(2) Equation (49) does not change the complex structure of the torus. Hence, only the information of the untransformed quantities $n_{\rm q}$ and $n_{\rm b}$ is needed to analyze the complex structure of the torus.
@@ V.1 Vanishing of the torus
In particular, it seems that Eq. (57) holds exactly when $\theta=0,\pi$ in LQCD simulation. See, e.g., Fig. 7 in Ref Sakai et al. (2009). and references therein. The phase of the modulus $\tau$ is given by
@@
At $\theta=\pi$, $n_{\rm b}$ is antiparallel to $n_{\rm q}$, and, hence, it is not linearly independent with $n_{\rm q}$. There, the torus shrinks smoothly to a one-dimensional object. Although there is no discontinuity of $\Delta\phi_{\mathrm{b}-q}$ below $T_{\rm RW}$, in this sense, the point $\theta=\pi$ is special. Of course, $n_{\rm b}$ is parallel to $n_{\rm q}$ also at $\theta=0$, but this fact is trivial since $\theta$ itself vanishes.
@@ V.2 Calculation of moduli in the fundamental region
Figure 4 shows the moduli $\tau$ and $\tau_{\rm f}$ when $T<T_{\rm RW}$ and $|\tau|=0.5$. $\tau$ and $\tau_{\rm f}$ are symmetric with respect to the line ${\rm Re}(\tau)=0$. There is a left-right symmetry. Although the torus vanishes at $\theta=\pi$, $\tau$ is continuous there.
Fig. 4: $\tau$ and $\tau_{\rm f}$ when $T<T_{\rm RW}$ and $|\tau|=0.5$. The dot-dashed line and the dots are $\tau$ and $\tau_{\rm f}$, respectively. $\tau(\theta\to 0)=(0.5,0)$, $\tau(\theta\to\pi\pm 0)=(-0.5,0)$ and $\tau(\theta\to 2\pi)=(0.5,0)$, while the absolute values of $\tau_{\rm f}(\theta\to 0)$, $\tau_{\rm f}(\theta\to\pi\pm 0)$ and $\tau_{\rm f}(\theta\to 2\pi)$ are large.
Fig. 5: $\tau$ and $\tau_{\rm f}$ when $T<T_{\rm RW}$ and $|\tau|=0.95$. The dot-dashed line and the dots are $\tau$ and $\tau_{\rm f}$, respectively. $\tau(\theta\to 0)=(0.95,0)$, $\tau(\theta\to\pi\pm 0)=(-0.95,0)$ and $\tau(\theta\to 2\pi)=(0.95,0)$ while the absolute values of $\tau_{\rm f}(\theta\to 0)$, $\tau_{\rm f}(\theta\to\pi\pm 0)$ and $\tau_{\rm f}(\theta\to 2\pi)$ are large.
@@
Fig. 8: $\tau$ and $\tau_{\rm f}$ when $T>T_{\rm RW}$ and $|\tau|=1.05$. The dot-dashed line and the dots are $\tau$ and $\tau_{\rm f}$, respectively. $\tau(\theta\to 0)=(1.05,0)$, $\tau(\theta\to\pi-0)=(-1.05/2,1.05\sqrt{3}/2)$, $\tau(\theta\to\pi+0)=(-1.05/2,-1.05\sqrt{3}/2)$, $\tau(\theta\to 2\pi)=(1.05,0)$, $\tau_{\rm f}(\theta\to\pi-0)=(0.475,0.909)$ and $\tau_{\rm f}(\theta\to\pi+0)=(0.475,-0.909)$ while the absolute values of $\tau_{\rm f}(\theta\to 0)$ and $\tau_{\rm f}(\theta\to 2\pi)$ are large.
Fig. 9: $\tau$ and $\tau_{\rm f}$ when $T>T_{\rm RW}$ and $|\tau|=2$. The dot-dashed line and the dots are $\tau$ and $\tau_{\rm f}$, respectively. $\tau(\theta\to 0)=(2.0,0)$, $\tau(\theta\to\pi-0)=(-1.0,\sqrt{3})$, $\tau(\theta\to\pi+0)=(-1.0,-\sqrt{3})$, $\tau_{\rm f}(\theta\to\pi-0)=(0,\sqrt{3})$ and $\tau_{\rm f}(\theta\to\pi+0)=(0,-\sqrt{3})$ while the absolute values of $\tau_{\rm f}(\theta\to 0)$ and $\tau_{\rm f}(\theta\to 2\pi)$ are large.
@@
Figure 7 shows the moduli $\tau$ and $\tau_{\rm f}$ when $T\to T_{\rm RW}+0$ and $|\tau|=1$. $\tau$ is not symmetric with respect to the line ${\rm Re}(\tau)=0$, but $\tau_{\rm f}$ is, since the left boundary of the fundamental region can be transformed into the right one by the T transformation and vice versa. Also note that $\tau_{\rm f}$ in Fig. 7 resembles that in Fig. 6, but the value of $\tau_{\rm f}(\theta\to\pi)$ is different. In this sense, the symmetry is broken even for $\tau_{\rm f}$ in Fig. 7. Also note that $\tau=(-0.5,\pm\sqrt{3}/2)$ is a fixed point of the modular transformation ${\cal{ST}}$. This transformation generates a discrete subgroup of the modular group, namely, ${\mathbb{Z}}_{3}$ group, since $({\cal{ST}})^{3}$ is an identity transformation. It is known that this property is related to the residual ${\mathbb{Z}}_{3}$-symmetry in modular flavor symmetry Kikuchi et al. (2022).
@@
In Fig. 10, we summarize the image of the RW transition based on the torus deformations. When $T<T_{\rm RW}$, the torus shrinks to the one-dimensional object at $\theta=\pi$ smoothly, and then the object transforms into another torus $\mathrm{Im}(\tau)$ of which has an opposite sign, but discontinuity does not appear. Since the modular transformation cannot change the sign of $\mathrm{Im}(\tau)$, the torus with a positive $\mathrm{Im}(\tau)$ cannot be smoothly transformed into the one with a negative $\mathrm{Im}(\tau)$ without shrinking of the torus. When, $T>T_{\rm RW}$, the torus changes discontinuously at $\theta=\pi$. The structure of the torus with modulus $\tau$ at high $T$ is similar to that with modulus $-1/\tau$ at low $T$ but the left-right symmetry is broken at high $T$.
@@
Since $s^{2}-1/4$ is real positive constant which does not depend on $\theta$, $\tau_{\mathrm{h}}$ forms a circle. The circle vanishes in the limit $|\tau|\to 0$. The center of the circle is given by